Draw the graph of the function f(n) = 1 - n
2
from the set of integers to set of integers.
1
Expert's answer
2020-09-08T15:13:19-0400

The function "f(n)=1-n^2, n\in\Z" is not one-to-one:

"f(-1)=1-(-1)^2=0"

"f(1)=1-(1)^2=0"

Hence "f(-1)=0=f(1)," but "(-1\not=1)"

The function "f(n)=1-n^2, n\in\Z" is not onto:

there is no integer "n" with "f(n)=1-n^2=2," for instance.

If "n\in\Z," then "-n\in\Z" and


"f(-n)=1-(-n)^2=1-n^2=f(n), n\in\Z"

The function "f(n)=1-n^2, n\in\Z" is even.